A NaCl crystal has a mass of 0.15mg. how many moles of Na^+ are present in the crystal. ( express the answer to two significant figures and include the appropriate units) . How many Na^+ ions are present? ( express the answer using two significant figures)
Answer #1

First Step: Convert 0.15 mg into grams. 0.15 mg / 1000 mg = 0.00015 g NaCl.


Second Step: convert it into moles. 0.00015 g NaCl / 58.44 g NaCl = 2.5*10-6 moles NaCl ( By using formulae: no of moles= mass/molar mass)  


Third Step: convert it into Na+ ions by setting conversion factor: 2.5*10-6 moles NaCl x (1 mol Na+ ions / 1 mol NaCl) = 2.5*10^-6 moles Na+ ions.

Fourth Step: for the Na+ ions present, all you have to do is convert it using avogadro number. So, set up conversion factor 2.5*10-6 moles Na+ ions x (6.022*1023 Na+ ions / 1 mol Na+ ions) = 1.50*1018 Na+ ions present.